"X=y\/bc-\u221aqr"


Find the value of r when b=2,c=1,q=3,x=2 and y= 4


1
Expert's answer
2021-07-19T05:51:03-0400
"\sqrt{qr}=\dfrac{y}{bc}-x"

"r=\dfrac{(\dfrac{y}{bc}-x)^2}{q}"

Substitute


"r=\dfrac{(\dfrac{4}{2(1)}-2)^2}{3}"

"r=0"